OpenStudy (anonymous):

um how do you prove that \(0=1\)?

OpenStudy (anonymous):

you dont because its not true lol

OpenStudy (anonymous):

show us the proof

TheSmartOne (thesmartone):

But it's not possible for this question because http://prntscr.com/6ij8ti

OpenStudy (anonymous):

i know one way but like any other ways? the way i know 0=1-1+1-1+1-1+1-1(repeat) 0=(1-1)+(1-1)... (repeat 0=1+(1-1)+(1-1)... repeat (1-1)=0 0=1+0+0+0+0....(repeat) 0=1?

OpenStudy (anonymous):

maybe it was a mistake in the proof

OpenStudy (anonymous):

maybe you might want to take this question to a philosophy teacher, where they teach with out logic

OpenStudy (anonymous):

lol well i was just curious if anyone like know this thing

TheSmartOne (thesmartone):

You are subtracting 1 from the 1 to make it 0. You can't do that without doing it to both sides.

OpenStudy (anonymous):

i know one way but like any other ways? the way i know 0=1-1+1-1+1-1+1-1(repeat) 0=(1-1)+(1-1)... (repeat 0=1+(-1+1)+(-1+1)... repeat (-1+1)=0 0=1+0+0+0+0....(repeat) 0=1?

OpenStudy (anonymous):

there sorry it was a mistake

OpenStudy (anonymous):

see... i used the power of a n infinity

OpenStudy (anonymous):

the repeat thing

TheSmartOne (thesmartone):

Well basically \(\sf 0\neq1\)

OpenStudy (anonymous):

it was just a hypothesis since if 0=1 then 0 could = everything... then we can get an 100 on our test or an 0 depending on what the teacher thinks.

TheSmartOne (thesmartone):

Haha

OpenStudy (anonymous):

ikr...

OpenStudy (anonymous):

then it would be scary since the hypothesis in the movie LUCY would be true and a pencil could turn into a universe...